

********************************************************************
UARS WEEKLY STATUS REPORT - 3 Jun 2005
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM remains powered off.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004273 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004274 and 2005039 has been processed and is waiting for archival
at the DAAC.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004259. MEPS data between 2004260 and
2005029 has been processed and is waiting for DAAC archival. MEPS data
for 2005030 to 2005039 is currently being processed.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered since updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 21 May 2005 (Orbit 74940, GMT Day 141) through
27 May 2005 (Orbit 75045, GMT Day 147).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 145/2249 75027 POWER OFF
The beta angle ranged between the angles of 31.8 degrees to 08.9
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 30 May 2005 (DOY 150).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 5.11 to 4.45 degrees C. Additional
battery performance data for each battery on 20 May 2005 and 27 May
2005 are provided in Appendix D.
The clock error ranged from 10.8 to -9.9 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 05
HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 02
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to closely monitor the status of the
batteries. The batteries maintained their VT-5 charge level until the
147/1948 event when level 6 was commanded. This was done in
preparation for the low beta/yaw period. The MPS was again able to
support HALOE throughout its turn on duration for this beta cycle as
well as day-only SOLSTICE operations.
Testing of the solar array drive resumed this week on 23 and 26 May
2005. The goal of the testing was to free the B side of the drive, so
that the A side could be utilized. Two attempts were made each day to
run through a series of slews to try to free the B side, but all
attempts were unsuccessful. The FOT will analyze the results of these
tests and try again next week.
MEETINGS
-None this report period.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 29 May 2005 (DOY 149). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 27 May 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
74982 143/1754 05-169 SADDE SADD TESTING
74983 143/1957 05-170 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
75002 145/0213 05-171 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
75015 145/2249 05-172 HALOE POWER OFF
75027 146/1744 05-173 SADDE SOLAR ARRAY TESTING
75043 147/1948 05-174 MPS VT LEVEL 6 (YAW)
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 27 May 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 27 May 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
74983 143/2003 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
75002 145/0225 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 27 May 2005
20 May 2005 - GMT Day 140 Beta = 30.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 11.2/ -05.6 -3.05/-4.63 +00.0/+00.0 05.8 0.00 0.0
2 257.6/ -56.0 +9.72/+8.39 +33.6/-20.8 25.1 1.04 24.4
3 201.6/-700.0 +2.91/+2.28 +18.0/-13.2 25.1 1.05 12.6
27 May 2005 - GMT Day 147 Beta = 08.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 16.8/ -05.6 -4.31/-5.26 +00.0/+00.0 05.8 0.00 0.0
2 11.2/-123.2 +7.08/+5.79 +33.2/-18.8 25.3 1.03 23.5
3 72.8/-700.0 +1.97/+1.34 +19.6/-12.0 25.3 1.05 12.6
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
75043 147/1948 LV5 -> LV6
APPENDIX E
UARS ANOMALY REPORTS
Week ending 27 May 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-029 FOT 74949 141/1423 39569 NO ODMS
05-028 FOT 74973 143/0438 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-028 FOT See List May 05 30832 Generic Late Acquisition
05-029 FOT 74949 141/1423 39569 NO ODMS